# Multi-Excitation Projective Simulation with a Many-Body Physics Inspired Inductive Bias

### v2

@@ I Introduction
Deep learning has become a powerful numerical tool, with various applications all over science and technology [1, 2, 3]. At the heart of this technological revolution are Artificial Neural Networks (ANN), parameterized function ansätze trained via gradient descent methods to achieve an ideal input-output behavior on data [4].
Despite the enormous success of ANNs, they also tend to have significant problems. First of all, their statistical nature means that ANNs will sometimes make mistakes, which can have dangerous consequences in high-risk applications such as medical diagnosis. This problem is amplified by the fact that ANNs are quite unreliable under so-called out-of-distribution data [5, 6, 7]: Data sampled from a different probability distribution that still has the same qualitative features which the ANN is supposed to learn. This weakness enables many adversarial attacks [8, 9]. Most importantly, the complex and mostly problem-agnostic structure of ANNs makes it difficult to understand their “reasoning process”, essentially turning ANNs into oracles. All these issues led to the development of the field known as eXplainable Artificial Intelligence (XAI) [10, 11].
One promising approach to XAI realizes that many conscious human decision-making processes take the form of a chain-of-thought. The most famous machine learning approach modelling chains-of-thought is in the setting of Large Language Models (LLM)[12, 13]. The framework of Projective Simulation (PS) [14, 15] combines a model of deliberation, based on episodic memory, with reinforcement learning  [16]. It thereby extracts the essential components of chains-of-thought, and realizes that deliberation processes can be understood as a random walk of a single particle on a graph with vertices representing concepts or thoughts. Since its first proposal in [14], PS has been successfully applied to many domains [17, 18, 19, 20, 21]. In most of these applications, vertices in the graph (referred to as clips) have a more basic interpretation, e.g., as remembered percepts, or actions, or sensorimotoric memories more generally. In this paper, we extend the interpretation of clips to ‘concepts’ and ‘thoughts’, to align with concurrent literature in XAI (but without claiming that clips convey the full meaning of these terms in a philosophical sense). Given this extension, the representation of chains-of-thought as simple paths in a graph is limited and cannot easily capture thoughts which are most naturally understood by taking their composite structure into account.
@@
The paper is organized as follows: First, in Section II, we describe Single-Excitation PS. In Section III, we present and define our Multi-Excitation PS agent, along with a dynamic hypergraph to model the agent’s training history in Subsection III. Then, in Subsection IV.1, we explain the physical motivation behind our inductive bias and later formalize it in Subsection IV.2. With this formalization, we derive complexity estimates in Subsection IV.3. The numerical experiments from the three learning scenarios we consider can be found in Section V. We then propose approaches towards an actual quantum mePS agent in Section VI. Finally, in Section VII, we discuss our results and suggest some promising future directions for the mePS framework.
@@
and initialized to 00. The glow dampening factor $\eta\in[0,1]$ is a hyperparameter, and plays a role very similar to the discount factor in returns and value functions [39]. The standard PS update rule can be interpreted as a form of Hebb’s learning rule “What fires together wires together”.
@@ III Multi-Excitation PS
For the purpose of interpretability, it is important to explicitly represent different observables and degrees of freedom. To make this possible, we first reimagine the random walk of PS as an excitation or a particle moving along the ECM. We will use the terms particle and excitation interchangeably. Now, to be capable of explicitly representing different observables as separate entities, we replace the single excitation with multiple excitations. With this change, it is also possible to have one clip for each value of each observable.
Figure 2: An example of a directed, weighted hypergraph describing the ECM of a typical mePS agent in a reinforcement learning setting. Percept clips are represented in blue with a lowercase $s$, intermediate clips in grey with a lowercase $c$, and action clips in red with a lowercase $a$; the domains and codomains of hyperedges are labelled with capital letters whose clip type corresponds to their lowercase version. Each hyperedge $e\in E$ also has an h-value $h(e)$ associated with it.
@@
A standard Multi-Excitation Projective Simulation (mePS) agent is given by a weighted, directed hypergraph $G=(V,E,h)$ that we refer to as the Episodic and Compositional Memory(ECM) of the agent (compare Figure 2). We refer to the elements $c\in V$ as clips and use the notation $V=\{c_{1},\ldots,c_{N}\}$. Subsets $C\subset V$ will be referred to as excitation configurations. Furthermore, we will often use the short-hand notation $c_{j}\equiv j$, identifying clips with their labels.
@@
[…]
Upon receiving an observation $\mathrm{obs}$, excitations are put on the (percept) clips in $\mathcal{I}(\mathrm{obs})$ in the agent’s ECM, triggering deliberation through the ECM (see Def. 5) until reaching a set of clips $C_{\mathrm{act}}$ contained in $\mathrm{OUT}$. Then, the action $\mathcal{O}(C_{\mathrm{act}})$ is used by the agent on the environment.
@@
Similarly to $h_{\mathrm{init}}$ and $g$, one may also consider introducing separate $\gamma$ and $\eta$ for each hyperedge; this of course comes at the cost of an explosion in the number of (hyper-)parameters, which is not a desirable feature of a good model.
@@
Let $T\subset\mathbb{R}$ be a parameter space with elements $t\in T$. Consider a weighted, directed hypergraph $G=(V,E,h)$ such that the vertex and hyperedge sets can be partitioned as $V=\bigcup_{t\in T}V_{t}$ and $E=\bigcup_{t\in T}E_{t}$, with $E_{t}\subset(\mathcal{P}(V_{t})\setminus{\varnothing})\times(\mathcal{P}(V_{t})\setminus{\varnothing})$, respectively. A weighted, directed, dynamic hypergraph $\mathcal{G}$ is a hypergraph foliation of sub-hypergraphs $\{G_{t}\}_{t\in T}$ where each $G_{t}=(V_{t},E_{t},h_{t})$ is called a leaf of the foliation. Each leaf has a weight function $h_{t}:E_{t}\rightarrow\mathbb{R}$ which corresponds to a domain restriction of the weight function $h:E\rightarrow\mathbb{R}$.
[…]
The above definition borrows the foliation concept from differential geometry that is often used to formulate initial value problems on Lorentzian manifolds in the theory of general relativity  [41]. The foliation structure in our definition allows us to explicitly relate all sub-hypergraphs appearing in the set through the identification of each of their weight functions $h_{t}$ as constant $t$ slices of the weight function $h$. In this way, the weight function $h$ acts as a sort of glue that stitches the sub-hypergraphs together, inducing a flow through the set. This is in stark contrast to the definition in [40] or other related definitions in the (hyper)graph visualization literature (to our knowledge) [42].
If we endow $h$ with the explicit form given in Definition 6, then the entire mePS algorithm can also essentially be viewed as a hypergraph generation tool, where hypergraphs with specific properties could be obtained after training by tailoring the agent architecture and update rule along with the learning environment. This process would produce a final hypergraph but if one also stores the generated hypergraphs at each training step, then the mePS algorithm can also generate dynamic hypergraphs, which could subsequently be analyzed using standard (hyper)graph visualization techniques [42, 43].
Because mePS is an explainable model, the dynamic hypergraph inherits this explainability and one can also visualize how the meaning of the sub-hypergraphs evolves over time. We believe the latter is an interesting application of (hyper)graph visualization to machine-learning training histories [44]. However, much infrastructure that is currently underdeveloped in normal PS implementations, such as the single-excitation PS graph surgery rules proposed in [14], needs to first be developed before such a proposal could be fruitfully initiated.
@@ IV A Physics-Inspired Inductive Bias
In this section, we will present our inductive bias that will allow us to significantly reduce the computational complexity of mePS agents. In Subsection IV.1, we motivate this inductive bias using inspiration from many-body physics. We put particular care into making this subsection digestible for non-physicists. However, readers not interested in the physical background can directly jump to Subsection IV.2 where we propose the formal specification of the inductive bias; this subsection does not require knowledge from Subsection IV.1. At last, in Subsection IV.3, we prove complexity bounds that establish the exponential reduction of computational complexity provided by our inductive bias. Furthermore, we prove additional reductions in computational complexity provided by modifications of the inductive bias adapted to layered hypergraphs.
IV.1 Motivation
So far, we discussed how to parametrize and train random walks involving several excitations in principle. In practice, however, one will face severe complexity issues if the random walks are allowed to traverse the entire hypergraph unrestricted. Each element of the power set $\mathcal{P}(V)$ of the clip set $V$ corresponds to a configuration of excitations, and there are $2^{|V|}$-many of such configurations. To tackle this problematic scaling, we propose an inductive bias motivated by quantum many-body physics. In physics, (quasi-) particles play the role of our excitations. Many if not most physical phenomena can be understood via models that only involve fundamental interactions of very few particles/excitations [32, 33]. In fact, in the standard model of particle physics [34, 35, 36], our most fundamental description of matter so far, there are no interactions involving more than four particles.
These observations motivate us to also think in terms of interactions of excitations and to introduce a cutoff on how many excitations can interact. For the remainder of this subsection, we will construct a classical analogue of the quantum dynamics of many-body systems to formalize our inductive bias. We emphasize again that we do not do any quantum machine learning in this work, only classical. A complete quantum mePS model will be the subject of a future work.
Each clip in our hypergraph can be interpreted as a mode in quantum many-body physics and in the formalism called second quantization, each excitation configuration is represented by a vector |$\ket{n_{1},\dots,n_{|V|}}$ in a Hilbert space, where $n_{j}$ is the number of excitations in clip/mode $j$.
Time evolution over the time duration $\Delta t$ is described by an operator $U(\Delta t)=e^{-i\hbar\Delta t\cdot H}$, where $i$ is the complex unit, $\hbar$ is the reduced Planck’s constant, $H$ is an operator called the Hamiltonian, and $e^{\bullet}$ is the matrix exponential. This means after a duration $\Delta t$, a many-body system starting in a state |$\ket{n_{1},\dots,n_{|V|}}$ is afterwards described by the state $e^{i\Delta t\cdot H}\ket{n_{1},\dots,n_{|V|}}$.
Important for our inductive bias are the typical expressions for $H$. For this, we require the ladder operators $a_{j}$ and $a^{\dagger}_{j}$, with ${}^{\dagger}$ denoting the hermitian adjoint (i.e. $A^{\dagger}=\overline{A^{T}}$, with $\overline{\bullet}$ denoting complex conjugation) and $j\in V$. $a^{\dagger}_{j}$ adds one excitation to clip $j$, i.e. $a^{\dagger}_{j}\ket{\ldots,n_{j},\ldots}\propto\ket{\ldots,n_{j}+1,\ldots}$ and is called a creation operator. Similarly, $a_{j}$ removes an excitation from a clip $j$, i.e. $a_{j}\ket{\ldots,n_{j},\ldots}\propto\ket{\ldots,n_{j}-1,\ldots}$, and is called an annihilation operator. For the special case $n_{j}=0$, we have $a_{j}\ket{\ldots,n_{j},\ldots}=0$.
A typical Hamiltonian $H$ in second quantization is of the form
For small enough time intervals $\delta t$, the time evolution operator (with $\hbar$ set equal to 1) can be approximated as $e^{iH\delta t}\approx\mathbb{1}+iH\delta t$. We discard the identity operator $\mathbb{1}$ (in physical terms, we post-select on a non-trivial change occurring) because it means that no transition occurred at all. We discretize time in multiples of $\delta t$, and identify each step in the random walk with one application of $i\delta tH$, absorbing $i\delta t$ into the coefficients $h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})$.
[…]
The last ingredient in our inductive bias is the observation that the operators $h(j_{1},\ldots,j_{i},k_{1},\ldots,k_{o})a^{\dagger}_{k_{1}}\cdots a^{\dagger}_{k_{o}}\cdot a_{j_{1}}\cdots a_{j_{i}}$ also act on excitation configurations |$\ket{n_{1},\ldots,n_{|V|}}$ that have excitations $n_{x}=1$ for $x\neq j_{1},\ldots,j_{i},k_{1},\ldots,k_{o}$, but leave $n_{x}$ unchanged. Similarly, we also apply $h$-values $h_{(i,o)}(\{c_{j_{1}},\ldots,c_{j_{i}}\},\{c_{k_{1}},\ldots,c_{k_{o}}\})$ to excitation configurations that have additional excitations in unrelated clips $c_{x}$.
IV.2 Formalization
Combining all of our considerations from subsection IV.1, our proposed inductive bias works as follows:
@@ IV.2 Formalization
    For all $(i,o)\in IO$, let $E^{(i,o)}_{\mathrm{all}}$ be the set of all $(C_{\mathrm{in}},C_{\mathrm{out}})\in(\mathcal{P}(V)\setminus\{\varnothing\})\times(\mathcal{P}(V)\setminus\{\varnothing\})$ with $|C_{\mathrm{in}}|=i$ and $|C_{\mathrm{out}}|=o$, and $C_{\mathrm{in}}\neq C_{\mathrm{out}}$. Here, the last condition serves to rule out transitions that do nothing. Then, specify a subset $E^{(i,o)}\subset E^{(i,o)}_{\mathrm{all}}$ which serves to describe the set of allowed transitions for $(i,o)$. The notation $C_{\mathrm{in}}\rightarrow C_{\mathrm{out}}$ will also be used for $e=(C_{\mathrm{in}},C_{\mathrm{out}})\in E^{(i,o)}$.
@@
    In the original configuration $\{c_{m_{1}},\ldots c_{m_{x}}\}$, remove all excitations in $\{c_{j_{1}},\ldots,c_{j_{i}}\}$, and put excitations into $\{c_{k_{1}},\ldots,c_{k_{o}}\}$. If ($\big{(}\{c_{m_{1}},\ldots,c_{m_{x}}\}\setminus\{c_{j_{1}},\ldots,c_{j_{i}}\}\big{)}\cap\{c_{k_{1}},\ldots,c_{k_{o}}\}\neq\varnothing$, we keep those excitations and discard the second excitations for those clips (see Remark 8).
Consider now the $n$-th step in the episode, and write an explicit time label ${}^{(n)}$ on the $h$-values and glows. Upon receiving a reward $R^{(n)}$, the many-body mePS update rule updates the many-body h-values $h^{(n)}_{(i,o)}:=h^{(n)}_{(i,o)}\big{(}C^{(i)}_{\mathrm{in}},C^{(o)}_{\mathrm{out}}\big{)}$ for all $(i,o)\in IO$ and all $C^{(i)}_{\mathrm{in}}\rightarrow C^{(o)}_{\mathrm{out}}\in E^{(i,o)}$ according to the rule
@@
It can happen that sampled transitions put excitations into clips that are already occupied. In Inductive Bias 1, we made the choice that the clip simply stays excited, i.e. it continues to carry exactly one excitation, effectively discarding the second excitation.
We made this choice because we associate clips with concepts or beliefs, and the excitation tells us whether the concept represented by the clip is currently relevant.
However, as we will explain in more detail in Section VI, this choice cannot naturally be linked to the behavior of any elementary particle. In fact, it is an irreversible process: the excitation that jumps on an already excited clip cannot jump back and is thereby annihilated.
We have assumed that there are only single shared hyperparameters $\gamma$ and $\eta$ for all transitions. Of course, one can modify the inductive bias and pick these hyperparameters separately for all transitions.
@@
Now, we have two hypergraphs, the ECM and the many-body hypergraph. Similarly, we have the standard $h$-values $h$ and the many-body $h$-values $h_{(i,o)}$. While conceptually related, it is not obvious that the definitions in Inductive Bias 1 are compatible with Definition 2. In Appendix B, we show that the definitions are indeed compatible during inference when using the standard probability assignment, by showing how to construct the standard $h$-values $h$ from the many-body $h$-values $h_{(i,o)}$.
Furthermore, it is important to emphasize that $h$ and $h_{(i,o)}$ are only consistent for inference, NOT during learning. Updating $h_{(i,o)}$ will also affect $h$ for transitions that did not occur in the random walk. However, for the rest of the paper, it is enough to only work with the $h_{(i,o)}$.
@@
Inductive Bias 2.
For weighted, layered feed-forward many-body hypergraphs with layers $(L_{1},\dots,L_{D})$, we introduce the following three modifications of Inductive Bias 1:
  1. FF.
@@
  2. SF.
    Let $\ell:V\rightarrow\{1,2,\dots D\}$ be the function that maps each clip to the layer it is in. Then, given an excitation configuration $\{c_{m_{1}},\dots,c_{m_{x}}\}$, with the labelling such that the layers satisfy $\ell(c_{m_{1}})=\cdots=\ell(c_{m_{n}})<\ell(c_{m_{n+1}})\leq\cdots\leq\ell(c_{m_{x}})$, the relevant $h_{(i,o)}(C_{\mathrm{in}},C_{\mathrm{out}})$-values for this configuration are restricted to those with $C_{\mathrm{in}}\subset\{c_{m_{1}},\dots,c_{m_{n}}\}$.
  3. DP.
@@
Also for these inductive biases, we will see in Appendix B that they are compatible with the standard mePS agent in Definition 2 during inference with the standard probability assignment.
Inductive Bias 2FF just enforces a feed-forward condition, and is the weakest of these Inductive Biases. In Section IV.3, we will see that it still allows for decision-making processes that take an exponential amount of time in the number of layers. Reducing the maximal deliberation time is the purpose of Inductive Bias 2SF and 2DP. Inductive Bias 2SF is the next stronger one. In Section IV.3, we will see that it reduces the maximal deliberation time to essentially the total number of clips.
However, there is one important consequence to keep in mind when modelling layered ECMs: In human decision-making, a common theme is to write down some intermediate results, and only use them much later when they are deemed relevant. An example would be the derivation of several independent lemmas, all of which get used in proving a theorem. Since in Inductive Bias 2SF the shallowest excitations are removed first, one should introduce copies of their clips in deeper layers to not lose the knowledge/concepts they represent in later steps.
Physically, this inductive bias corresponds to situations encountered for example in integrated photonics chips performing quantum computation with several photons: The photons have a fixed lateral velocity in the interferometer circuit, but perform a quantum walk in the transversal direction [47, 48]. A common noise source of such chips is that photons get absorbed by the environment.

### v3

# Multi-Excitation Projective Simulation with a Many-Body Physics-Inspired Inductive Bias

@@ I Introduction
Deep learning has become a powerful numerical tool, with various applications all over science and technology [1, 2, 3]. At the heart of this technological revolution are Artificial Neural Networks (ANN), parameterized function ansätze trained via gradient descent methods to achieve an ideal input-output behavior on data [4]. Despite the enormous success of ANNs, their complex and mostly problem-agnostic structure makes it difficult to understand their “reasoning process”, essentially turning ANNs into oracles. This, along with other pertinent issues related to reliability and trustworthiness [5, 6, 7, 8, 9], led to the development of the field known as eXplainable Artificial Intelligence (XAI) [10, 11].
One promising approach to XAI realizes that many conscious human decision-making processes take the form of a chain-of-thought. The most famous machine learning approach modelling chains-of-thought is in the setting of Large Language Models (LLM)[12, 13]. The framework of Projective Simulation (PS) [14, 15] combines a model of deliberation, based on episodic memory, with reinforcement learning  [16]. It thereby extracts the essential components of chains-of-thought, and realizes that deliberation processes can be understood as a random walk of a single particle on a graph with vertices representing concepts or thoughts. Since its first proposal in [14], PS has been successfully applied to many domains [17, 18, 19, 20, 21]. In most of these applications, vertices in the graph (referred to as clips) have a more basic interpretation, e.g., as remembered percepts, or actions, or sensorimotoric memories more generally. However, the representation of chains-of-thought as simple paths in a graph is limited and cannot easily capture thoughts which are most naturally understood by taking their composite structure into account.
@@
Since the inductive bias still allows for cyclic hypergraphs, the random walk has an infinite worst-case running time. This motivates us to formulate modifications of our inductive bias and the corresponding hypergraphs for which low order polynomial runtimes are guaranteed. More specifically, we prove worst case upper bounds for the runtimes of the random walks for layered hypergraphs. These worst case paths are rarely encountered in practice, so we leave the formulation and formal analysis of these inductive biases for the appendix.
The paper is organized as follows: First, in Section II, we describe Single-Excitation PS. In Section III, we present and define our Multi-Excitation PS agent, along with a dynamic hypergraph to model the agent’s training history in Subsection III. In Section IV, we develop the formalization of our inductive bias, including the exponential reduction in trainable parameters. However, since the motivation involves the quantum physical description of many-body systems, we will delay the presentation of this motivation towards the end in Section VI.Before, in Section V, the numerical experiments from the three learning scenarios we consider can be found in Section V. Afterwards, as already stated, we present the quantum many-body physical motivation of our classical inductive bias in Section VI.1. Based on this quantum physical foundation, we then propose approaches towards an actual quantum mePS agent in Section VI.2. Finally, in Section VII, we discuss our results and suggest some promising future directions for the mePS framework.
In Appendix A, we present more details on the numerical experiments. In Appendix B, we present modifications of our inductive bias for layered hypergraphs which guarantee polynomial upper bounds on the number of steps in the random walk. Afterwards, in Appendix C, we prove that our inductive bias agents are fully compatible with the definition of mePS agents given in Section III. Then, in Appendix D, we prove the aforementioned bounds for different variants of the inductive bias.
@@
and initialized to 00. The glow dampening factor $\eta\in[0,1]$ is a hyperparameter, and plays a role similar to the discount factor in returns and value functions [39]. The standard PS update rule can be interpreted as a form of Hebb’s learning rule “What fires together wires together”.
@@ III Multi-Excitation PS
For the purpose of interpretability, it is important to explicitly represent different observables and degrees of freedom. To make this possible, we first imagine the random walk of PS as an excitation or a particle moving along the ECM. We will use the terms particle and excitation interchangeably. Now, to be capable of explicitly representing different observables as separate entities, we replace the single excitation with multiple excitations. With this change, it is also possible to have one excitation for each value of each observable.
Figure 2: An example of a directed, weighted hypergraph describing the ECM of a typical mePS agent in a reinforcement learning setting. Atomic percept clips are represented in blue with a lowercase $s$, atomic intermediate clips in grey with a lowercase $c$, and atomic action clips in red with a lowercase $a$; the domains and codomains of hyperedges are labelled with capital letters whose clip type corresponds to their lowercase version. Each hyperedge $e\in E$ also has an h-value $h(e)$ associated with it.
@@
A standard Multi-Excitation Projective Simulation (mePS) agent is given by a weighted, directed hypergraph $G=(V,E,h)$ that we refer to as the Episodic and Compositional Memory(ECM) of the agent (compare Figure 2). We refer to the elements $c\in V$ as atomic clips and use the notation $V=\{c_{1},\ldots,c_{N}\}$. Subsets $C\subset V$ will be referred to as excitation configurations. Furthermore, we will often use the short-hand notation $c_{j}\equiv j$, identifying atomic clips with their labels.
@@
Let $\mathrm{OUT}\subset P(V)\setminus\{\varnothing\}$ and $\mathrm{IN}\subset P(V)\setminus\{\varnothing\}$ denote output and input sets, respectively. In the setting of Markov Decision Processes, a mePS agent is also equipped with the following two functions: an input coupling function $\mathcal{I}:\text{Observations}\rightarrow\mathrm{IN}$ and an output coupling function $\mathcal{O}:\mathrm{OUT}\rightarrow\text{Actions}$.
Upon receiving an observation $\mathrm{obs}$, excitations are put on the atomic (percept) clips in $\mathcal{I}(\mathrm{obs})$ in the agent’s ECM, triggering deliberation through the ECM (see Def. 5) until reaching a set of atomic clips $C_{\mathrm{act}}$ contained in $\mathrm{OUT}$. Then, the action $\mathcal{O}(C_{\mathrm{act}})$ is used by the agent on the environment.
@@
    We refer to this list as the relevant $h$-values for the next random walk step.
@@
Let $T\subset\mathbb{R}$ be a parameter space with elements $t\in T$. Consider a weighted, directed hypergraph $G=(V,E,h)$ such that the vertex and hyperedge sets can be partitioned as $V=\bigcup_{t\in T}V_{t}$ and $E=\bigcup_{t\in T}E_{t}$, with $E_{t}\subset(\mathcal{P}(V_{t})\setminus{\varnothing})\times(\mathcal{P}(V_{t})\setminus{\varnothing})$, respectively. A weighted, directed, dynamic hypergraph $\mathcal{G}$ is a collection of sub-hypergraphs $\{G_{t}\}_{t\in T}$ where each $G_{t}=(V_{t},E_{t},h_{t})$ is called a leaf of $\mathcal{G}$. Each leaf has a weight function $h_{t}:E_{t}\rightarrow\mathbb{R}$ which corresponds to a domain restriction of the weight function $h:E\rightarrow\mathbb{R}$.
The initialization $h_{\text{init}}$ in particular is the weight function corresponding to the smallest leaf index $\min_{t\in T}t$, assuming a minimal index exists.
Our definition allows us to explicitly relate all sub-hypergraphs appearing in the set through the identification of each of their weight functions $h_{t}$ as constant $t$ slices of the weight function $h$. In this way, the weight function $h$ acts as a sort of glue that stitches the sub-hypergraphs together, inducing a flow through the set. This is in stark contrast to the definition in [40] or other related definitions in the (hyper)graph visualization literature (to our knowledge) [41].
If we endow $h$ with the explicit form given in Definition 6, then the entire mePS algorithm can also be viewed as a hypergraph generation tool, where hypergraphs with specific properties could be obtained after training by tailoring the agent architecture and update rule along with the learning environment. This process would produce a final hypergraph but if one also stores the generated hypergraphs at each training step, then the mePS algorithm can also generate dynamic hypergraphs, which could subsequently be analyzed using standard (hyper)graph visualization techniques [41, 42]. Much infrastructure that is currently underdeveloped in normal PS implementations, such as the single-excitation PS graph surgery rules proposed in [14], would need to first be developed before such a proposal could be fruitfully initiated. However, the inductive bias that will be described later in Section IV may provide a temporary solution to the problem of generating a hypergraph with an arbitrary size, essentially by adaptively restricting the size of the subsets appearing in the edge set $E$ during training.
Because mePS is an explainable model, the dynamic hypergraph inherits this explainability and one can also visualize how the meaning of the sub-hypergraphs evolves over time. We believe the latter is an interesting application of (hyper)graph visualization to machine-learning training histories and XAI in general [43].
@@ IV A Physics-Inspired Inductive Bias
In this section, we will present our inductive bias that will allow us to significantly reduce the computational complexity of mePS agents. Since the detailed motivation is inspired by quantum many-body physics to formulate classical analogues, we postpone a detailed explanation of the motivation to Section VI.1. Nonetheless, we emphasize that we have written said section in an accessible way which we believe to be digestible also for non-physicists.
However, already here, we provide a basic, conceptual discussion motivating the main decisions for our inductive bias. The main cause for the exponential complexity of the hypergraph is that we assign transition probabilities or $h$-values $h(C_{\mathrm{in}},C_{\mathrm{out}})$ for full excitation configurations $C_{\mathrm{in}/\mathrm{out}}$, in a hypergraph that can have exponentially many such configurations.
In many-body physics, one also investigates transition probabilities of many excitations. Such transition probabilities can usually be understood as arising from combinations of elementary interactions of only a handful of excitations. In general, these interactions can destroy or create excitations, a typical example being nuclear scattering.
This motivates us to introduce a finite set $IO$ which contains tuples $(i,o)$ of integers. For each $(i,o)\in IO$, there exists an elementary transition which converts $i$ ingoing excitations into $o$ outgoing excitations. Typically in the most well-known and successful physical models (e.g. the Ising model) $i,o\leq 4$, so that only a small number of excitations are required to effectively capture a wide range of phenomena. In our classical analogy, we measure the “strength” or amplitude of this elementary transition with $h$-values $h_{(i,o)}(C^{(i)}_{\mathrm{in}},C^{(o)}_{\mathrm{out}})$. This amplitude depends not only on the number of scattering excitations, but also on the precise state of the $i$ ingoing excitations $C^{(i)}_{\mathrm{in}}$ and the $o$ outgoing excitations $C^{(o)}_{\mathrm{out}}$. In our inductive bias, these many-body $h$-values $h_{(i,o)}(C^{(i)}_{\mathrm{in}},C^{(o)}_{\mathrm{out}})$ will be the trainable parameters.
Since these $h_{(i,o)}$ express elementary interactions/transitions, we allow for the presence of other excitations. This means that $C^{(i)}_{\mathrm{in}}\subset C_{\mathrm{in}}$ and $C^{(o)}_{\mathrm{out}}\subset C_{\mathrm{out}}$, i.e. the ingoing excitations and outgoing excitations of an elementary interaction are only subsets of the full excitation configurations now and later. However, physics does not allow arbitary transitions between particles. Conservation laws such as charge conservation and locality restrict the allowed transitions. We use sets $E^{(i,o)}\subset\mathcal{P}(V)\times\mathcal{P}(V)$ to define which elementary transitions we allow, i.e. only for $(C^{(i)}_{\mathrm{in}},C^{(o)}_{\mathrm{out}})\in E^{(i,o)}$ we assign (trainable) $h_{(i,o)}$.
In our classical analogy, we make the convention that one random walk step corresponds to one elementary transition occurring. Furthermore, as already stated, we measure the strength for each elementary process with the $h_{(i,o)}$. Therefore, to sample the next excitation configuration, we just make a list of all the $h_{(i,o)}(C_{\mathrm{in}}^{(i)},C_{\mathrm{out}}^{(o)})$ such that $C_{\mathrm{in}}^{(i)}\subset C_{\mathrm{in}}$, and directly sample a $C_{\mathrm{out}}^{(o)}$ from these $h_{(i,o)}$, using e.g. the standard probability assignment or a softmax. Then, the next full excitation configuration is simply given by $C_{\mathrm{out}}:=C^{(o)}_{\mathrm{out}}\cup(C_{\mathrm{in}}\setminus C^{(i)}_{\mathrm{out}})$.
Now, after this exposition, we give the full formal definition of our inductive bias, including its update rule:
@@
    For all $(i,o)\in IO$, let $E^{(i,o)}_{\mathrm{all}}$ be the set of all $(C^{(i)}_{\mathrm{in}},C^{(o)}_{\mathrm{out}})\in(\mathcal{P}(V)\setminus\{\varnothing\})\times(\mathcal{P}(V)\setminus\{\varnothing\})$ with $|C^{(i)}_{\mathrm{in}}|=i$ and $|C^{(o)}_{\mathrm{out}}|=o$, and $C^{(i)}_{\mathrm{in}}\neq C^{(o)}_{\mathrm{out}}$. Here, the last condition serves to rule out transitions that do nothing. Then, specify a subset $E^{(i,o)}\subset E^{(i,o)}_{\mathrm{all}}$ which serves to describe the set of allowed transitions for $(i,o)$. The notation $C^{(i)}_{\mathrm{in}}\rightarrow C^{(o)}_{\mathrm{out}}$ will also be used for $e=(C^{(i)}_{\mathrm{in}},C^{(o)}_{\mathrm{out}})\in E^{(i,o)}$.
@@
    In the original configuration $\{c_{m_{1}},\ldots c_{m_{x}}\}$, remove all excitations in $\{c_{j_{1}},\ldots,c_{j_{i}}\}$, and put excitations into $\{c_{k_{1}},\ldots,c_{k_{o}}\}$. If ($\big{(}\{c_{m_{1}},\ldots,c_{m_{x}}\}\setminus\{c_{j_{1}},\ldots,c_{j_{i}}\}\big{)}\cap\{c_{k_{1}},\ldots,c_{k_{o}}\}\neq\varnothing$, we keep those excitations and discard the second excitations for those atomic clips (see Remark 8).
Consider now the $n$-th step in the episode, and write an explicit time label (n) on the $h$-values and glows. Upon receiving a reward $R^{(n)}$, the many-body mePS update rule updates the many-body h-values $h^{(n)}_{(i,o)}:=h^{(n)}_{(i,o)}\big{(}C^{(i)}_{\mathrm{in}},C^{(o)}_{\mathrm{out}}\big{)}$ for all $(i,o)\in IO$ and all $C^{(i)}_{\mathrm{in}}\rightarrow C^{(o)}_{\mathrm{out}}\in E^{(i,o)}$ according to the rule
@@
It can happen that sampled transitions put excitations into atomic clips that are already occupied. In Inductive Bias 1, we made the choice that the atomic clip simply stays excited, i.e. it continues to carry exactly one excitation, effectively discarding the second excitation.
We made this choice because we associate atomic clips with concepts or beliefs, and the excitation tells us whether the concept represented by the atomic clip is currently relevant.
However, as we will explain in more detail in Section VI, this choice cannot naturally be linked to the behavior of any elementary particle. In fact, it is an irreversible process: the excitation that jumps on an already excited atomic clip cannot jump back and is thereby annihilated.
@@
Now, we have two hypergraphs, the ECM and the many-body hypergraph. Similarly, we have the standard $h$-values $h$ and the many-body $h$-values $h_{(i,o)}$. While conceptually related, it is not obvious that the definitions in Inductive Bias 1 are compatible with Definition 2. In Appendix C, we show that the definitions are indeed compatible during inference when using the standard probability assignment, by showing how to construct the standard $h$-values $h$ from the many-body $h$-values $h_{(i,o)}$.
Furthermore, it is important to emphasize that $h$ and $h_{(i,o)}$ are only equivalent for inference, NOT during learning. Updating $h_{(i,o)}$ will also affect $h$ for transitions that did not occur in the random walk. However, for the rest of the paper, it is enough to only work with the $h_{(i,o)}$.
@@
Inductive Bias 2FF.
For weighted, layered feed-forward many-body hypergraphs with layers $(L_{1},\dots,L_{D})$, we introduce the following modification of Inductive Bias 1:
@@
We are referring to this Inductive Bias with the label 2FF because in Appendix B we will introduce some small modifications of this inductive bias called 2SF and 2DP. These modifications guarantee a polynomial upper bound on the random walk path lengths. However, in the main text, we will focus on Inductive Bias 2FF because the worst-case paths are rarely encountered in practice.
Also for this inductive bias, we will see in Appendix C that it is compatible with the standard mePS agent in Definition 2 during inference with the standard probability assignment.
Now that we have formulated our Inductive Bias and its main variant for layered ECMs, we provide an example that discusses how these inductive biases get applied.
Figure 3: An example illustrating random walk steps under different inductive biases, compare with Example 11. Excited atomic clips are shown in red. The sampled hyperedge is shown in blue. Subfigure a) shows a deliberation step which is only allowed under Inductive Bias 1, because its codomain is in two layers. Also, it shows that an excitation moving into an occupied atomic clip gets discarded. Subfigure b) shows a typical transition under Inductive Bias 2FF.